Please show all work and provide explanation for the following:
(i) You, have two sets of data注刈MTOV want to id out the relationship diiedl hon of tese vauiables. X is a sample of SLoves on a tesl (per student out of loo) and Y i6 the line ㅓhi Same Sudent spent Studying Gn mnules). Round your answer ta two a as after each calculation 1.1JEnd 扌he median, mode, and range. 4 oeach data se 貂 (a) Fnd the coefficient of correlation 8e 47 of these two data sels ㅓken interpret fo his number 78 44 33 4 3Plet dkese poinis on a scatter 90 48 diaaram 3S 47 38 ② Four candidates , are, running for mayor, Adam Brown, Collins, and Dalton assuming hay all haue an equal chance of election mne sample space what if we imd out the probabliies of eledion are assigned respective厶: a relatile frequency 9abei answer ㅔe follouing lna is the pcbabiy that Adams loses? that Brown wins?
Answer #1

x

y

x^2

y^2

xy

77

40

5929

1600

3080

63

42

3969

1764

2646

79

37

6241

1369

2923

86

47

7396

2209

4042

51

25

2601

625

1275

78

44

6084

1936

3432

83

41

6889

1681

3403

90

48

8100

2304

4320

65

35

4225

1225

2275

47

28

2209

784

1316

Ex

Ey

Ex^2

Ey^2

Exy

719

387

53643

15497

28712

Mean=Ex/n

Mean (x)=719/10=71.9

Mean (y)=387/10=38.7

Median=(n+1)/2 nd value

Median(x)=77.5

Median(y)=40.5

Mode=Most frequently repeated value

Mode(x)=N/A

Mode(y)= N/A

correlation: -

n(Exy)-(Ex)(Ey)/sqrtnEx^2-(Ex)^2* nEy^2-(Ey)^2

10*28712-(719)(387)/ sqrt10*53643-(719)^2* 10*15497-(387)^2

0.88

50 30 20 10 100 80 60 40 20